OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -3 2x - 6y = -6 A. -1 B. zero C. 1 D. 2

OpenStudy (anonymous):

Take the second equation: 2x -6y = -6 add 6y to both sides: 2x = -6 +6y divide both sides by 2: x = -3 +3y Take other equation: 5x - 3y = -3 Substitute x equation: 5(-3+3y) - 3y = -3 which is: -15 +15y - 3y = -3 now add 15 to both sides, combine terms and reduce

OpenStudy (anonymous):

Thanks!